Test the series:


n=1 ∑∞ (-1)n-1 [Sin(nx)]/n√n


for absolute and conditional convergence.


1
Expert's answer
2021-06-16T06:25:55-0400
"\displaystyle\sum_{n=1}^{\infin}\dfrac{(-1)^n\sin[nx]}{n\sqrt{n}}"

"|\sin[nx]|\leq1, x\in\R, n\geq1"

Then


"\bigg|\dfrac{(-1)^n\sin[nx]}{n\sqrt{n}}\bigg|\leq\dfrac{1}{n\sqrt{n}}, x\in\R, n\geq1"

The series "\displaystyle\sum_{n=1}^{\infin}\dfrac{1}{n\sqrt{n}}" converges as "p" -series with "p=\dfrac{3}{2}>1."


Therefore the series "\displaystyle\sum_{n=1}^{\infin}\dfrac{(-1)^n\sin[nx]}{n\sqrt{n}}" converges absolutely by the Comparison Test, for "x\in\R."